AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:08 2025
Author
*
:
Subject
*
:
<p>We have recovered the beam at 01:49, after such a long time we had no target so we increased from 80K to 120K at 01:52. 02:01 update 120K-->95K as we have observed around that temperature always fluctuations and is safer.</p> <p>Ring current 7.74e7 particles/s</p> <p><strong>Merger, Tape server (WS3): </strong>1619 kB/s writing ok</p> <p><strong>Leakage current (WS4): 3.26 uA </strong></p> <p><strong>DAQ (WS2): temperature checked, clock status: ok</strong><strong>, adc calibration: ok, white rabbit check: failed</strong></p> <p><strong>Online monitoring (WS 5): done</strong></p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown